You have a trust that mature to $75,000 that you will receive when you are 40...

  1. You have a trust that mature to $75,000 that you will receive when you are 40 years old. You are 23 years old now. If the trust is growing at 4% per year what is the value of the trust now? ____________
Answer #1

Let the value of the trust today = P

I takes 17 (40 - 23) years to become $75,000. Growth rate = 4% per year

P * (1 + 0.04)17 = $75,000

P * (1.04)17 = $75,000

P* 1.9479 = $75,000

P = $ 38,503

Value of the trust now is $38,503